OpenStudy (anonymous):

The table below shows the values of y for each value of x. x y 0 1 1 2 2 3 3 3 Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. (4 points) Part B: The function f(h) shown below represents a plumber's total charges, inclusive of call out charges, for h hours of work. f(h) = 70h + 50 Calculate and interpret the meaning of f(2). (4 points) Part C: Write an ordered pair to represent the input and output of the function in Part B when the plumber works for d

OpenStudy (ali2x2):

keystone?

OpenStudy (welshfella):

Part A It is a function as there are no duplicate values of x

OpenStudy (welshfella):

f(h) = 70h + 50 f(2) = 70(2) + 50 = ?

OpenStudy (welshfella):

50 is the call out charge and he charges 70 dollars for every hour worked.

OpenStudy (welshfella):

f(2) is the charge for 2 hours work.